### v1

# Searching for periodic signals and quasi-periodic oscillations from an extremely active cycle of FRB 20240114A

Abstract
[…]
radio transit source — radio bursts — time series analysis
facilities: FASTsoftware: numpy (Harris et al., 2020), matplotlib (Hunter, 2007), astropy (Astropy Collaboration et al., 2022), PINT (Luo et al., 2021; Susobhanan et al., 2024), emcee (Foreman-Mackey et al., 2013), riptide (Morello et al., 2020), presto (Ransom, 2011), stingray (Huppenkothen et al., 2019), lmfit (Newville et al., 2014), PyEMD (Laszuk, 2017)
1 Introduction
Fast radio bursts (FRBs), first discovered in 2007 (Lorimer et al., 2007), are millisecond-duration, highly energetic radio transients of extragalactic origin. These transient and energetic bursts exhibit dispersion measures (DMs) that significantly exceed the Galactic contribution, thereby strongly suggesting their cosmological distances and placing them among some of the most intriguing astrophysical phenomena (Lorimer et al., 2007; Thornton et al., 2013; CHIME/FRB Collaboration et al., 2021), though large host-galaxy DM contributions have also been found in some cases (e.g., FRB 20190520B; Niu et al. 2022a). Since their initial discovery, the field of FRB research has grown rapidly, with extensive observational campaigns leading to the identification of hundreds of these mysterious bursts. These detections have revealed a diverse population, with some FRBs appearing to repeat sporadically while others have only been observed as singular, non-repeating events (Petroff et al., 2019; CHIME/FRB Collaboration et al., 2021; Xu et al., 2023). The identification of repeating FRBs, most notably FRB 20121102A (Spitler et al., 2016), has ruled out the early speculations that all FRBs originate from cataclysmic one-time events. Instead, the existence of repeaters indicates that at least some FRBs are powered by non-catastrophic engines.
A breakthrough in the field of FRB research was marked by the detection of FRB 20200428, which was definitively associated with an X-ray burst from the Galactic magnetar SGR J1935+2154 (Bochenek et al., 2020; Mereghetti et al., 2020; Li et al., 2021a). Another targeted multi-wavelength campaign did not detect further coincident radio bursts from this source, suggesting that such FRB-SGR associations are rare and might require extreme physical conditions to produce coherent radio emission (Lin et al., 2020). Extensive observations of repeating FRBs have revealed non-Poissonian burst clustering, long-term variations in DM and rotation measure (RM), as well as complex waiting time statistics, highlighting their highly stochastic nature and posing challenges to single rotating compact object models (Li et al., 2021b; Hu & Huang, 2023; Zhang et al., 2024; Liu et al., 2025). Recent polarization and magnetospheric studies suggest that repeating FRBs reside in complex, evolving environments consistent with young magnetars, as indicated by rapid RM variations and strong circular polarization, although observations of FRB 20220912A, which shows high circular polarization but nearly constant RM, suggest that at least some repeaters may not be embedded in strongly magneto-ionic environments (Xu et al., 2022; Feng et al., 2022; Zhu et al., 2023; Feng et al., 2024, 2025). A successful search for pulsar-like, spin-modulated burst-to-burst temporal periodicities in FRB sources would provide a convincing observational evidence that at least some FRBs originate from compact stars.
Some FRB sources have demonstrated distinct long-timescale periodic activity. Notable examples include FRB 20180916B, which follows a 16-days cycle of activity (Chime/Frb Collaboration, 2020; Pastor-Marazuela et al., 2021; Sand et al., 2023), and FRB 20121102A, which has shown signs of a much longer periodicity of approximately 160 days, alongside a shorter periodicity of around 4.6 days (Rajwade et al., 2020; Cruces et al., 2021; Li et al., 2024; Braga et al., 2025). The recently discovered FRB 20240209A, located at the outskirts of a quiescent elliptical galaxy (z = 0.1384), exhibits a periodic activity cycle of about 126 days, further adding to the rare class of periodically repeating FRBs (Pal, 2025). Furthermore, FRB 20220529 shows a potential periodicity of approximately 200 days in its RM, which is consistent with a binary origin, as suggested by the significant RM variation and its rapid recovery (Liang et al., 2025). Similarly, FRB 20201124A exhibits a periodicity of 26 days in its RM, further supporting the idea that periodic RM variations may be a common feature of actively repeating FRBs in binary systems (Xu et al., 2025). The origin of periodic activity in repeating FRBs remains debated, with proposed explanations falling into several categories. Rotation-powered models suggest the periodicity reflects the spin of isolated magnetars (Xu et al., 2021; Lan et al., 2024). Binary models attribute the modulation to orbital dynamics in neutron star systems, including Be/X-ray binaries (Ioka & Zhang, 2020; Lyutikov et al., 2020; Zhang & Gao, 2020; Li et al., 2021c; Barkov & Popov, 2022). Precession effects have also been proposed as a possible cause of the observed periodicity (Yang & Zou, 2020; Tong et al., 2020; Levin et al., 2020; Zanazzi & Lai, 2020; Katz, 2020; Chen, 2020; Chen et al., 2021; Sridhar et al., 2021; Wei et al., 2022). Alternative explanations involve asteroid belt interactions (Dai & Zhong, 2020; Voisin et al., 2021; Deng et al., 2024) or the repetitive collapse of the crust of an accreting strange quark star (Geng et al., 2021). The diversity of models highlights the need for further observational constraints, as many fundamental questions about the physical origins and emission mechanisms of FRBs remain open (Zhang, 2023).
Short-timescale periodicity in FRBs provides critical constraints on their emission mechanisms and progenitors. Some bursts have shown millisecond-scale quasi-periodic structures, e.g. the 0.411 ms spacing in FRB 20201020A (Pastor-Marazuela et al., 2023), although its significance remains marginal. These sub-second periodicities challenge the spin explanations due to extreme rotational requirements, favoring magnetospheric processes instead. Additionally, the detection of ultra-narrow spike components in FRB 20230607A, down to 0.3 ms, further supports a magnetospheric origin for at least some repeating FRBs (Zhou et al., 2025). On the other hand, other studies have found no significant periodicity in repeating FRBs like FRB 20201124A, ruling out isolated magnetars or binaries with certain parameters (Niu et al., 2022b). Recent claims of a $\sim$1.7 s period in FRB 20201124A (Du et al., 2025), if confirmed, could indicate a pulsar origin, while low-significance periodicity candidates such as the 7.267 ms signal in FRB 20230708A (Dial et al., 2025) highlight the challenges in distinguishing true periodicity from burst microstructure. These controversial results highlight the need for higher-cadence and higher time resolution observations to clarify the nature of short-timescale FRB variability.
This study systematically analyzes the periodicity of FRB 20240114A, an extremely active FRB. The source was discovered by the Canadian Hydrogen Intensity Mapping Experiment (CHIME) in January 2024, and since then, it has been observed by multiple telescopes (Shin & CHIME/FRB Collaboration, 2024; Tian et al., 2024; Xie et al., 2024; Huang et al., 2025). The source has been continuously observed by the FRB key project of FAST since January 2024. The analysis is based on a dataset collected from January 28 to August 30, 2024, during which 11,553 burst events were recorded (Zhang Junshuo et al. 2025; Zhang Longxuan et al. 2025; Wang et al. 2025, all in preparation). This extensive sample offers the most comprehensive dataset to date for investigating the presence of multiple timescale periodicities in FRBs, aiding in the identification of periodic signals and QPOs from the source. This paper is organized as follows. In Section 2, we search for periodic signals in the TOAs data of FRB 20240114A to explore possible links to magnetar rotation, binary orbits, and precession scenarios. In Section 3, we mainly investigate QPOs in the burst time series, which provide crucial insights into the underlying radiation mechanisms. Section 4 discusses the implications of our findings, and Section 5 summarizes the main conclusions of this study.
Figure 1: The 57 observations of FAST on FRB 20240114A from January 28 to August 30, 2024, showing the distributions of burst counts (bars in the upper panel), burst rates (red dots in the upper panel), and exposure times (bars in the lower panel). Bar charts with the same color indicate that a consistent short-timescale periodic signal may exist in the corresponding observations (see Section 2).
2 Searching for Periodic Signals in TOA data
The TOAs in this paper are derived from 57 FAST observations of FRB 20240114A between January 28 and August 30, 2024. For more details on the data processing of this source, including burst searching, dispersion correction, and TOA definition, please refer to the paper by Junshuo Zhang et al. (2025, in preparation). The processed data, including the TOAs and other derived parameters, can be accessed via https://doi.org/10.57760/sciencedb.Fastro.00030. Since the data span several months, PINT (Luo et al., 2021; Susobhanan et al., 2024) was used to convert the TOAs to the solar system barycenter. Figure 1 shows the time distribution of bursts across the 57 observations. We perform periodic signal searches on the TOA data at both short and long timescales.
2.1 Method
2.1.1 Short-Timescale Periodicity Search
Short-timescale searches refer to conducting periodic searches on the obtained TOA data, covering timescales ranging from several tens of milliseconds to tens of minutes. A dual analysis approach was employed for each continuous observation dataset. This involved first performing a wide-range scan of frequency $f$ and its first derivative $\dot{f}$ using the Time-Differencing Algorithm (TDA), followed by precise period determination through the $Z^{2}$ statistic.
Table 1: Candidate periodic signals detected by the TDA and refined using the $Z_{n}^{2}$ statistic. The table lists the candidate number, observation date (MJD), parameters from the TDA detection (frequency $f$, derivative $\dot{f}$, period $P$, derivative $\dot{P}$, significance $\sigma$), and the refined parameters from the $Z_{n}^{2}$ statistic.
No. MJD TDA Detection $Z_{n}^{2}$ Refinement
$f$ $\dot{f}$ $P$ $\dot{P}$ $\sigma$ $f$ $\dot{f}$ $P$ $\dot{P}$
Hz ×$\times 10^{-7}$ Hz/s s ×$\times 10^{-7}$ s/s Hz ×$\times 10^{-7}$ Hz/s s ×$\times 10^{-7}$ s/s
1 60401 1.4868 -14.556 0.6726 6.585 4.9 1.4869 -16.301 0.6726 7.374
60403 1.4848 0.000 0.6735 0.000 4.1 1.4854 -8.310 0.6732 3.767
2 60398 1.5759 5.478 0.6346 -2.206 3.4 1.5761 2.304 0.6345 -0.927
60519 1.5739 7.770 0.6354 -3.137 6.0 1.5735 10.384 0.6355 -4.194
3 60417 1.8648 -2.534 0.5363 0.729 3.7 1.8653 -9.304 0.5361 2.674
60517 1.8665 8.259 0.5358 -2.371 3.5 1.8664 8.923 0.5358 -2.561
The TDA has been widely applied in periodicity detection for high-energy gamma-ray pulsars (Atwood et al., 2006; Ziegler et al., 2008; Abdo et. al., 2008; Abdo et al., 2010; Smith et al., 2023). This method is particularly suitable for sparse event signals like FRBs that exhibit only a few burst events per day. The TDA evaluates periodicity by analyzing time differences between events across a sliding time window. The statistical quantity searched by this method is (Atwood et al., 2006)
$D_{l}^{N_{w}}=\sum_{m=k+1}^{M}\sum_{k=0}^{N-1}a_{k}a_{m}e^{-i2\pi l(m-k)/N},$ (1)
where $a_{k}$ represents the number of TOAs in the kth time sample and $N$ is the number of time samples in the overall viewing period $T_{v}$. The upper limit for the outer summation is given by $M={\rm min}(N-1,k+N/N_{w})$, where $N/N_{w}=NT_{\rm win}/T_{v}$ represents the number of time samples within the window defined by $T_{\rm win}$. When $|{\rm Re}(D_{l}^{N_{w}})|$ exceeds a predetermined threshold, the corresponding frequency is identified as a candidate.
Our search simultaneously covers both frequency $f$ and its derivative $\dot{f}$, accounting not only for possible spin-down effects of the FRB source but also for potential binary system scenarios, provided that $T_{v}\leq 10P_{\rm b}$, where $P_{\rm b}$ is the orbital period of the FRB source (Ransom et al., 2003). The correction for $\dot{f}$ can be determined through
$t_{i}=\tilde{t_{i}}+\frac{1}{2}\frac{\dot{f}}{f}\tilde{t_{i}}^{2},$ (2)
where $\tilde{t_{i}}$ and $t_{i}$ represent the original and corrected TOAs respectively. In our search, the frequency range extends from the lower bound determined by the observation duration up to 32 Hz, with two harmonics combined to enhance the signal-to-noise ratio. The step size for searching $\frac{\dot{f}}{f}$ is given by
$d\left(\frac{\dot{f}}{f}\right)=\frac{1}{T_{\rm win}T_{v}f_{\rm max}},$ (3)
where $f_{\rm max}$ represents the maximum searched frequency. This parameter space covers plausible physical scenarios for FRB sources. $T_{\rm win}$ balances computational efficiency and frequency resolution: for observations $<$ 1800 s, $T_{\rm win}=T_{v}$; for observations $\geq$ 1800 s, $T_{\rm win}$ is fixed at 1800 s. Previous studies by Wang et al. (2022) suggest that repeating FRBs might originate from magnetars in binary systems with Be stars possessing disk structures. The range of $\dot{f}/f$ explored spans physically plausible regimes for both spin-down and binary motion. In a circular binary orbit, this quantity relates to the orbital period $P_{\rm b}$ and companion mass $M_{\rm c}$ via
$\left|\frac{\dot{f}}{f}\right|_{\rm max}=\frac{a_{\rm max}}{c}=\frac{(2\pi)^{\frac{4}{3}}G^{\frac{1}{3}}M_{c}}{(M_{p}+M_{c})^{\frac{2}{3}}}P_{b}^{-\frac{4}{3}}c^{-1}.$ (4)
Following Niu et al. (2022b), we adopt an upper limit of $M_{c}=100M_{\odot}$ to represent an extreme massive companion. Under this assumption, $\left(\dot{f}/f\right)_{\max}=10^{-6}\ \mathrm{s^{-1}}$, which corresponds to an orbital period of $P_{b}\sim 8.2$ hr. We also consider potential spin-down effects of the FRB source. For magnetars, the magnetic field strength of the surface $B_{s}$ is related to the spin period $P$ and its derivative $\dot{P}$ through $B_{s}=3.2\times 10^{19}\,\mathrm{G}\sqrt{P\dot{P}}$. Assuming a magnetar surface field of $10^{15}\,\mathrm{G}$, the intrinsic frequency evolution yields $\frac{\dot{f}}{f}\sim-9.77\times 10^{-10}f^{2}\ \mathrm{s^{-1}}$. Considering the known spin frequency of magnetars [^1] https://www.physics.mcgill.ca/ pulsar/magnetar/main, the value of $\frac{\dot{f}}{f}$ ranges from $-9.5\times 10^{-9}\ \mathrm{s^{-1}}$ to $-7.13\times 10^{-12}\ \mathrm{s^{-1}}$.
Our search strategy includes three complementary approaches:
  1. 1.
    Dedicated search: The TOAs from each single continuous observation are analyzed separately, covering frequency $f$ and its derivative $\dot{f}$ with $\frac{\dot{f}}{f}$ ranging from $-10^{-6}$ to $10^{-6}\,\mathrm{s}^{-1}$.
  2. 2.
    Joint search across days: All TOAs data from multiple observations are searched collectively, with a refined $\frac{\dot{f}}{f}$ range of $-10^{-8}$ to $0\,\mathrm{s}^{-1}$ to focus on long-term spin-down trends.
  3. 3.
    Clustered search: Since a single $\dot{f}$ may not fully capture the frequency evolution over long timescales, the TOAs are divided into segments (each $\leq$ 3 days) and analyzed independently. For each segment, $\frac{\dot{f}}{f}$ is constrained to $-10^{-8}$ to $0\,\mathrm{s}^{-1}$, allowing for a more precise characterization of the spin-down behavior within shorter intervals.
The search range is designed to account for both long-term spin-down evolution and potential binary orbital dynamics, thereby ensuring sensitivity to strongly magnetized magnetars in either scenario.
[…]
$Z^{2}_{n}=\frac{2}{N}\sum_{k=1}^{n}\left[\left(\sum_{j=1}^{N}\cos k\phi_{j}\right)^{2}+\left(\sum_{j=1}^{N}\sin k\phi_{j}\right)^{2}\right],$ (5)
where N is the number of TOAs and $\phi_{j}=2\pi(ft_{j}+\frac{1}{2}\dot{f}t_{j}^{2})$. We evaluate the significance of each candidate by computing $Z^{2}_{n}$ over a fine grid around the TDA-derived $(f,\dot{f})$, incorporating the first $n=4$ harmonics to balance sensitivity and computational cost. This two-stage approach, which combines a wide search using the TDA with subsequent $Z^{2}_{n}$ refinement, ensures efficient and reliable detection of weak periodic signals in sparse TOAs.
2.1.2 Long-Timescale Periodicity Search
The long-timescale search aims to detect periodicity on timescales of the order of days, which corresponds to the active period of the FRB. Two complementary techniques are employed in this study: epoch folding with $\chi^{2}$ testing and Lomb-Scargle periodogram (LSP) analysis.
Figure 2: Folding results of the corresponding TOAs, based on the parameters in Table 1. From top to bottom: profiles of the three candidate signals and their corresponding time-phase plots.
The epoch folding search spans $P_{\rm min}=1.5$ d to $P_{\rm max}=T/3$ d, where $T=214.46$ d is the time difference between the first and last TOA. The frequency step size is $df=0.1/T$, which determines the corresponding period step size through the relation between $P$ and $f$. We compute weighted mean TOA for each observation, with weights proportional to the daily burst rate. The $\chi^{2}$ statistic evaluates the significance of periodicity:
$\chi^{2}=\sum_{i=1}^{n}\frac{(N_{i}-\langle N\rangle_{i})^{2}}{\langle N\rangle_{i}},$ (6)
where $N_{i}$ represents the weighted TOA counts in phase bin $i$, $\langle N\rangle_{i}=aT_{i}$ is the expected count under the null hypothesis of constant rate, with $T_{i}$ being the exposure time and $a$ the mean burst rate. In our periodicity analysis, we systematically tested phase bin configurations ranging from 4 to 16 bins. The candidate selection process incorporated several rigorous criteria: the first of which was that only signals exceeding the $3\sigma$ confidence level (with appropriate multiple testing corrections) were to be considered significant. When detecting multiple adjacent period candidates above the threshold, we merged these and retained solely the most statistically significant period. To ensure physical validity, the candidates were required to appear consistently in at least two independent bin configurations. Furthermore, we implemented a sidereal day rejection filter to exclude any spurious signals matching integer multiples ($N=1$-$10$) of the characteristic 0.99727-day sidereal period within a specified tolerance window.
Figure 3: Long-timescale periodicity search results for the TOA data. Panel a: Epoch folding search result, with green dash-dotted and red dashed lines representing the $3\sigma$ and $5\sigma$ thresholds, respectively. The blue solid line indicates the degrees of freedom. The red solid line denotes a signal exceeding the $3\sigma$ threshold around 69.18 days. Panel b: Similar to panel a, with the difference that it uses the Lomb-Scargle algorithm, detecting a $3\sigma$ excess signal around 67.51 days. Panel c: Profile obtained by folding the TOA data with a period of 69.18 days. The black bar and red dashed line indicate the profile with and without exposure time correction, respectively. Panel d: The relative exposure time distribution as a function of phase, calculated for the signal’s period.
For the LSP analysis, we use the weight of each observed TOA as the intensity and the average TOA from each observation as the time. We then conduct a periodicity search within the same range mentioned above. We use the False Alarm Probability (FAP) to quantify the detection significance. The FAP is defined as (VanderPlas, 2018)
${\rm FAP}(z)\approx 1-\left[P_{\rm single}(z)\right]^{N_{\rm eff}},$ (7)
where $P_{\rm single}(z)=(1-z)^{\frac{N-1}{2}}$ represents the cumulative probability of observing a peak less than z at a specific frequency, assuming Gaussian noise. $N_{\rm eff}=f_{\rm max}T$, where $T$ is the time span of the time series, and $f_{\rm max}$ is the maximum frequency of the search.
2.2 Results
2.2.1 Short-Timescale Periodicity
We first performed a joint search of all observational data, which did not yield any candidates meeting our selection criteria. Subsequently, we conducted an analysis of individual observations by grouping those taken within 2–3-day intervals. This clustering approach resulted in 14 distinct groups of TOA data being analyzed, none of which produced candidates that satisfied our criteria. The individual observation analysis revealed several candidates that met our selection criteria. We require each observation to include at least three TOAs to be included in the search, which resulted in the exclusion of one observation from MJD 60344 out of the total 57 observational datasets. We detected 2765 candidates exceeding the $3\sigma$ significance level after multiple testing correction. Given the wide search parameter space and sparse TOA distribution, such a number of apparent candidates is statistically expected. Therefore, we imposed stringent criteria to isolate only the most robust signals: cross-epoch recurrence and minimum $P$-value within frequency clusters. First, for candidates with similar $f$ (where similarity is defined as $\Delta f/f<10^{-3}$) yet different $\dot{f}$, only the candidate with the smallest $P$-value was retained. Subsequently, we retained candidates that reappeared in at least two separate observation epochs. This rigorous process ultimately identified three candidates exhibiting cross-epoch recurrence. The dates on which these signals appeared are marked in the same color in Figure 1.
For precise parameter determination, we performed a refined search in their $f$-$\dot{f}$ parameter space using the $Z_{n}^{2}$ statistic. Specifically, for each candidate signal, we searched for 16,000 combinations of $f$ and $\dot{f}$ centered on the values determined by the TDA, with step sizes chosen such that the maximum phase drift over the entire continuous observation period would be less than 0.01. The $Z_{n}^{2}$ analysis led to the identification of three distinct candidates: one with a period of $\sim$0.673 s (detected in observations 60401 and 60403), another with a period of $\sim$0.635 s (detected in observations 60398 and 60519), and a third with a period of $\sim$0.536 s (detected in observations 60417 and 60517). Table 1 presents the sifted candidates identified by the TDA through rigorous selection criteria, alongside their refined parameters obtained via precision analysis with the $Z_{n}^{2}$ statistic. Figure 2 displays the profile and the corresponding time-phase diagram using these optimized parameters.
2.2.2 Long-Timescale Periodicity
[…]
3 searching for (quasi-)periodic signals in burst time series data
3.1 Method
To detect periodic signals and QPOs in FRB time series, we employed two complementary analysis methods, focusing respectively on the structures within individual bursts and on the temporal patterns among burst clusters. These approaches differ primarily due to the distinct properties of the data. Burst time series typically exhibit complex non-stationary structures that resist decomposition into independent subcomponents. Consequently, frequency-domain analysis based on the Fast Fourier Transform (FFT) and Bayesian inference is required. Since the individual bursts within a burst cluster are relatively independent, measurable parameters such as subpulse widths and amplitudes can be utilized. These parameters, combined with the autocorrelation function (ACF) and phase randomization methods, facilitates the evaluation of the significance of periodic signals, thereby enabling the effective periodicity searches.
Figure 4: Four quasi-periodic candidates identified from the burst time series using the ACF method, each with a significance exceeding $3\sigma$. Each candidate consists of four subplots: the top-left subplot shows the burst time series, the bottom-left subplot shows the corresponding dynamic spectrum, the top-right subplot shows the ACF of the burst time series, and the bottom-right subplot shows the simulated statistic used to calculate the probability of generating this ACF pattern. The red line in the burst time series represents the periodic intervals determined by the ACF analysis, while the red line in the simulated statistic corresponds to the observed value.
For periodicity detection of bursts, we adopted the Bayesian framework proposed by Huppenkothen et al. (2013). The method begins by computing the periodogram of burst time series via the FFT, then fits a null hypothesis model to obtain a maximum a posteriori (MAP) parameters. We calculate residuals $R_{j}=2I_{j}/S_{j}$, where $I_{j}$ represents the observed periodogram and $S_{j}$ denotes the null model’s prediction. Markov Chain Monte Carlo (MCMC) sampling generates simulated periodograms, establishing a distribution of maximum residuals against which observed residuals are compared for significance assessment. In order to detect QPOs, we further introduce a nested alternative hypothesis model, computing the likelihood ratio test (LRT) statistic $T_{\mathrm{LRT}}^{\mathrm{obs}}$ and determining its significance through simulated periodograms.
For periodicity detection of burst clusters, we employed the ACF method proposed by Kramer et al. (2024). The method first requires determining an appropriate time window. By analyzing the distribution of burst waiting times, we found that the distribution exhibits a bimodal structure with a transition around 0.2 seconds. Based on this, bursts with intervals shorter than 0.2 seconds are clustered into the same time series. Periodic signals in ACF analysis manifest as equally spaced local maxima. We quantify the strength of periodicity using the statistic $\mathcal{P}=\sum\rho(P_{\mu}\cdot i)$, where $\rho$ denotes the power of the ACF, and $P_{\mu}\cdot i$ represents integer multiples of the period $P_{\mu}$. The original subpulse counts, widths, and amplitudes are preserved, while uniformly distributed perturbations are incorporated at periodic intervals. The simulated sequences then undergo the same ACF analysis to generate a distribution for comparison with the observed values, thereby obtaining the $P$-value.
The periodicity detection method for individual bursts circumvents the challenges of decomposing complex signals into subcomponents. Conversely, the method for burst clusters leverages the timing characteristics of discrete events to facilitate the assessment of periodic significance. For details of the methodology, see references Chime/Frb Collaboration et al. (2022); Kramer et al. (2024) for the ACF-based analysis and Appendix A for the implementation of the FFT-based approach.
Figure 5: Example of a QPO embedded in a burst time series (No. 1 in Table LABEL:tab:QPO_pars). A1: Periodogram of the time series before DM removal. A2: Power spectrum of the time series after DM removal with a DM of 528.5 $\rm pc/cm^{3}$. Two peaks can be clearly seen around 260Hz and 500Hz. The fitting results using three Gaussian models (one fitting the low-frequency broadband component and two fitting the QPO) and a constant model (fitting white noise) are shown. A3: Variation curve of the relative intensity of the QPO with a central frequency of around 260Hz with respect to different DMs. B1: Dynamic spectrum of the burst, with red short lines indicating the frequency channels where RFI has been removed. B2: Fitting results of the burst time series using an asymmetric Gaussian model and a constant model. B3: Residuals of the burst time series fit. C1: The QPO parameters obtained from the fit in A2. C2: Periodogram of the burst time series after the best fit in B2, showing that the QPO is still present. C3: Periodogram of the residuals shown in B3, with no significant QPO observed. Based on C2 and C3, it is clear that the QPO signal exists in the pulse profile, rather than in the low-amplitude oscillations of the residuals.
3.2 Results
We applied both the ACF and the FFT methods to detect periodic signals and QPOs in the burst time series.
For the ACF analysis, we clustered bursts using a 0.2-second time window, resulting in 516 time series. The ACF was calculated for each of these time series, and the minimum lag of the significant peaks was identified as the candidate period $P_{\mu}$. The significance of these periodicities was assessed using the methodology outlined in Section 3.1. We identified 4 candidates with significance greater than $3\sigma$. These burst time series, their dynamic spectra, ACFs, and distributions of simulated statistics are shown in Figure 4. These time series exhibit relatively few peaks, with those peaks spaced at regular intervals. Most of these candidates exhibit 3 peaks, with one candidate showing 4 peaks, resulting in a high significance of up to $3.9\sigma$. In Appendix C (Figure 10), we also list 7 additional candidates with significance lower than $3\sigma$ yet still showing some evidence of periodicity. Notably, some of these candidates match the period well up to 3 peaks, although the agreement breaks down after the 4th peak, leading to lower significance. This suggests that the candidates with significance greater than $3\sigma$ may not exhibit periodic patterns in longer time series. The periods of these candidates range from 5 ms to 62 ms, reflecting the clustering of bursts within a 0.2-second time window. As expected, increasing the time window length makes it more difficult to maintain periodicity. This aligns with the results from searching for short-timescale periods in the TOA data, where pulse phases exhibited considerable spread rather than remaining confined to a narrow range. The 0.2-second clustering assumption implies that bursts occurring within such short intervals are likely to be physically connected, indicating the potential for periodicity in these local bursts.
Figure 6: Posterior distributions of the null hypothesis model parameters and the likelihood ratio test. Upper panel: The corner plot displays the posterior distributions of the model parameters $\mu$, $\sigma$, and $c$ from the null hypothesis employed to fit the low-frequency component and white noise in the periodogram (see Figure 5A2). Here, $\mu$, $\sigma$, and $c$ represent the amplitude, centre value, width of the Gaussian function, and white noise level, respectively. Lower panel: A comparison of the observed likelihood ratio with the distribution of $2\times 10^{6}$ likelihood ratios obtained from the posterior distributions. The observed value significantly deviates from the sample distribution, indicating a rejection of the null hypothesis. For further details, refer to Section 3.2.
[…]
In summary, both the ACF and the FFT methods provided valuable insights into the periodic and quasi-periodic characteristics of the observed bursts. The ACF analysis revealed periodicity in short-timescale burst clustering, with characteristic timescales ranging from a few milliseconds to tens of milliseconds. In parallel, the FFT analysis identified QPOs with frequency components in the hundreds of Hz, thereby suggesting the presence of high-frequency oscillatory behavior. The complementary nature of these methods allows us to gain a comprehensive understanding of the periodic phenomena present in the burst time series.
4 Discussion
4.1 Short-Timescale Periodicity in TOAs
A natural expectation from magnetar-based models of repeating FRBs is the presence of underlying periodicities in burst arrival times that trace the rotation of the neutron star. Although a large-scale joint analysis revealed no significant global periodicity, consistent with the non-detection reported by Niu et al. (2022b), we conducted detailed searches within each continuous observation and identified three sets of candidate periodicities, each observed in two separate epochs.
Although we have identified three candidate periodic signals from FRB 20240114A, it is unlikely that all of them correspond to genuine periodicity to the source. Our current selection criteria do not permit us to definitively exclude any of the candidates. While the periods of these candidate signals are on the shorter side of the known magnetar period range, some known magnetars, such as PSR J1846-0258, have even shorter periods of $\sim$0.3266 s (Livingstone et al., 2011).
A striking feature of these candidate signals is the large day-to-day variation in their measured period derivatives ($\dot{P}$), including sign reversals. These variations are difficult to reconcile with the intrinsic spin-down expected from isolated magnetars. Instead, they are more naturally interpreted as apparent period derivatives caused by line-of-sight acceleration in a binary system. If FRB 20240114A resides in such a system, orbital motion could induce variable Doppler shifts, leading to fluctuations in both $P$ and $\dot{P}$ over timescales of days. This scenario is further supported by the fact that the magnitudes of the measured $\dot{P}$ values often exceed those typically associated with magnetar spin-down, sometimes by several orders of magnitude. The system has the potential to yield significant insights into binary dynamics and the role of companion stars in shaping FRB emission.
The absence of substantial periodic detections in our joint and clustered analyses suggests that any underlying periodicity is likely to be transient, intermittent, or strongly modulated, consistent with findings from other FRB studies (Pastor-Marazuela et al., 2023; Du et al., 2025). Du et al. (2025) propose that FRB production may involve multiple independent emission regions in the magnetar magnetosphere, potentially explaining this intermittency. Should these periodicities be confirmed, they would suggest the presence of extremely young, highly magnetized neutron stars as FRB progenitors. If the binary system hypothesis holds, it could also shed light on the role of companion stars in modulating the magnetar’s rotation. Future multi-wavelength observations will be crucial in verifying these periodicities and constraining the system’s properties, including the potential presence of a companion star.
4.2 Long-Timescale Periodicity in TOAs
In the activity of FRB 20240114A, a candidate periodicity of 69.18 days with a significance of 3.9 $\sigma$ was detected using the epoch folding method, while a similar candidate periodicity of 67.51 days with a significance of 3.5 $\sigma$ was detected using Lomb-Scargle analysis. These results suggest that this repeating FRB may exhibit periodic active behavior around a 68-day period. Concerns about the reliability of this candidate periodicity have been raised due to the uneven distribution of exposure time with respect to phase. Although the Lomb-Scargle periodogram is a useful tool for identifying periodic signals in unevenly sampled data, it is susceptible to observational biases, particularly when the exposure time distribution is uneven. This issue has also been observed in the analysis of other periodic FRBs, such as FRB 20180916B and FRB 20121102A, where the significance of the detected periodicity was influenced by the observation cadence and the exposure time distribution (Chime/Frb Collaboration, 2020; Rajwade et al., 2020).
If the candidate period of approximately 68 days is confirmed, FRB 20240114A would join a growing class of periodic repeating FRBs. The diversity in periods among these sources suggests that multiple possible mechanisms could explain their periodic behavior. For FRB 20180916B, various models have been proposed including neutron star spin periods modulated by disk-wind interactions (Xu et al., 2021), binary orbital modulation with either massive stellar companions or compact objects (Zhang & Gao, 2020; Ioka & Zhang, 2020; Li et al., 2021c), asteroid belt interactions (Dai & Zhong, 2020), and precession mechanisms ranging from free neutron star precession to jet precession in accreting systems (Zanazzi & Lai, 2020; Tong et al., 2020; Chen et al., 2021). Similarly, the $\sim$157-day periodicity of FRB 20121102A has been attributed to either orbital motion or precession effects (Ioka & Zhang, 2020; Levin et al., 2020). The 68-day period of FRB 20240114A, if confirmed, could represent an intermediate case between these established periodic FRBs, potentially favoring binary models with longer orbital periods or different companion types.
The tentative nature of FRB 20240114A’s periodicity highlights the challenges in identifying periodic signals with uneven observational coverage. Future observations should prioritize uniform cadence to minimize sampling biases and multiwavelength follow-up to constrain the local environment. The growing diversity of FRB periods suggests multiple physical mechanisms may be at work, requiring continued observational and theoretical efforts to unravel their origins.
4.3 Periodicity and QPOs in Burst Time Series
By combining the ACF and the FFT analysis, we reveal temporal features in FRB emissions that are both distinct and complementary. Using the ACF-based method, we identified 4 candidates that exceed the $3\sigma$ threshold. Among them, only one shows four distinct pulses with clearly periodic burst times, indicating a regular temporal spacing. The detected periodicities (ranging from 5 to 62 ms) in the burst clusters exhibit an intriguing behavior: their statistical significance diminishes when the analysis is extended beyond 4 cycles. This suggests a form of transient coherence, possibly originating from unstable oscillation modes in the emission region. These modes may temporarily excite specific periodic components, leading to enhancements in periodicity. The lack of persistent periodic signals over longer timescales is consistent with the stochastic nature of burst arrival times. At the same time, the localized synchronization observed within short windows implies that some degree of phase correlation exists during active emission episodes. Using the FFT analysis, we identified 5 candidates exhibiting statistically significant QPO features. Although the periodogram characteristics could suggest additional candidates, we adopted a conservative selection strategy, retaining only those with unambiguous oscillatory signatures visible in both the time-domain waveforms and their corresponding periodograms.
Figure 7: The result of decomposing the burst time series shown in Figure 5B2 using the EMD. The EMD yields 6 IMFs. Panel a: The original total burst time series. Panel b: The time series obtained by superimposing IMFs 1-3 from the EMD, corresponding to white noise. Panel c: The time series obtained by superimposing IMFs 4-5 from the EMD, corresponding to the QPO time-domain signal. Panel d: The time series obtained from IMF 6 from the EMD, corresponding to the external profile of the burst. Panels e through h: The periodograms corresponding to the time series in panels a through d. The pink curve in panel c represents the best fit of the QPO time-domain profile using model (8).
To further analyze the temporal characteristics of these QPOs, we applied Empirical Mode Decomposition (EMD) to decompose the burst time series into multiple scales. The EMD is an adaptive signal processing technique that decomposes complex nonlinear, non-stationary signals into Intrinsic Mode Functions (IMFs), each representing a specific time-frequency component (Huang et al., 1998; Barbosh et al., 2020). We used a Python implementation of the algorithm to achieve our goal (Laszuk, 2017). We use the QPO candidate shown in Fig. 5 as an example for the analysis. Through the EMD, we obtained 6 IMFs covering different frequency bands. Following the hierarchical transition from high to low frequencies in the IMFs, we combined high-frequency IMFs 1-3 (Component 1), mid-frequency IMFs 4-5 (Component 2), and low-frequency IMF 6 (Component 3), and then calculated periodograms for each component. Figure 7a-d respectively show the original time series and the time series of the three components obtained through the EMD, while Figure 7e-h display their corresponding periodograms. Component 1 primarily consists of high-frequency white noise, and Component 3 represents the envelope of the original burst time series. Component 2 exhibits oscillatory behavior in its time series, with a periodogram consistent with the QPOs in the original periodogram, confirming that Component 2 contains the QPOs extracted via the EMD.
We adopt the method used in Li et al. (2022) to fit the time-domain signals of QPOs, employing the product of a Gaussian function and a sine function to model the QPOs extracted by the EMD. This method is based on the assumption that QPOs are amplitude-modulated signals, and its mathematical expression is
$I(t)=Ae^{-\frac{(t-t_{c})^{2}}{2\sigma^{2}}}\sin(2\pi ft+\phi),$ (8)
where $A$ represents the maximum amplitude of the signal, $t_{c}$ is the center position of the Gaussian function, $\sigma$ is the width of the Gaussian function, $f$ is the oscillation frequency, and $\phi$ is the initial phase. Using this model to fit the data, we obtain the optimal parameters as $A=1.047$, $t_{c}=0.026$ s, $\sigma=0.003$ s , $f=270.380$ Hz, $\phi=-2.300$ rad. The oscillation frequency $f$ is consistent with the central frequency of the QPOs obtained by directly fitting the periodogram. This further verifies the presence of QPOs in the original burst time series. The corresponding burst time series, QPO time-domain signals extracted via the EMD, the fitting results using equation (8), and their corresponding periodograms are all presented in Appendix C (Figure 9).
Figure 8 displays the distribution characteristics of these QPOs’ central frequencies and full width at half maximum (FWHM). The results show that both the FWHM and the central frequencies of the QPOs exhibit unimodal distributions, with a significant correlation between them (Pearson correlation coefficient $R=0.97$). Linear regression yields a relationship of $y=0.32x+19.99$. This strong correlation suggests that the quality factor ($Q\equiv f_{c}/\mathrm{FWHM}$, where $f_{c}$ is the central frequency of the QPOs) remains relatively stable across different burst time series, indicating that the underlying physical mechanism may be universal.
The microstructure of FRBs provides crucial clues for understanding their radiation mechanisms and the physical properties of their central engines. Among these features, the discovery of QPOs is particularly intriguing. The presence of QPOs may be closely related to various physical processes in compact objects, such as rotation, gravitational lensing, binary star mergers, crustal oscillations in magnetars, or instabilities in the magnetosphere (Piro, 2012; Wang et al., 2016; Wadiasingh & Chirenti, 2020; Pastor-Marazuela et al., 2023). These periodic or quasi-periodic phenomena likely reflect the dynamic behavior of compact objects under extreme magnetic fields or the oscillation modes of magnetospheric plasma. In-depth studies of these QPO features hold the potential to offer new insights into the origin of FRB radiation and their connection to compact objects.
Figure 8: The joint distribution of QPO frequencies and their FWHM in FRB 20240114A, with marginal distributions shown alongside. To avoid harmonic interference, only the lowest frequency QPO from each burst time series is selected as a sample for the plot. The parameters show a strong positive correlation (Pearson $R=0.97$).
A possible physical scenario involves a periodic pulse train superimposed on a broader pulse envelope, which could arise from temporal modulation of the signal strength (e.g., through gravitational lensing effects in pulsars). This mechanism naturally produces a bell-shaped overall profile with short-timescale periodic sub-structures, and may also be considered in the context of FRB emission models. In contrast, the pulse signals we observed do not exhibit strict periodicity, instead showing QPOs. Therefore, the gravitational lensing model is even less effective in explaining the phenomenon we observed. They also explored the possibility of explaining the periodic signal through neutron star mergers. Neutron star mergers are expected to produce QPOs ranging from a few Hz to several kHz, although the merger process is extremely brief. The duration typically depends on the companion star’s mass, ranging from approximately $\sim 10^{2}-10^{4}$ s, or between 0.1 and 1 s (Chime/Frb Collaboration et al., 2022). In contrast, FRB 20240114A is a repeater, which has been erupting for over a year, with no signs suggesting that the bursts will end in the short term. Therefore, the QPOs of FRB 20240114A cannot originate from a neutron star merger.
The microstructure of FRB pulses, especially the appearance of QPOs, may be related to the magnetar’s X-ray burst activity and could be closely associated with the magnetar’s vibrations or intense magnetic field activity (Thompson et al., 2017; Suvorov & Kokkotas, 2019; Li et al., 2021a, 2022).
[…]
We detected QPOs only in a subset of radio pulses, likely because substructure detectability depends critically on time resolution and the signal-to-noise ratio, combined with possible source intensity variations. Even under similar observational conditions, quasi-periodic subpulses are not universally present. For example, the magnetar PSR J1622–4950 clearly exhibited such structures in 2017 observations, although they were relatively weak in observations from early and late 2018 (Kramer et al., 2024).
5 Summary
This study presents a comprehensive multi-timescale analysis of periodic signals and QPOs in the repeating fast radio burst FRB 20240114A, based on extensive observations with the FAST. Our investigation reveals several key findings across different temporal domains:
  1. 1.
    Short-Timescale Periodicity Search: We analyzed 11,553 bursts from 57 observations spanning January to August 2024. We identified three candidate periodic signals with periods of 0.673 s, 0.635 s, and 0.536 s (significance $>3\sigma$), each detected in two independent observations. It is clear that not all of these signals are related to the periodicity of the source. Nevertheless, our current selection criteria are not sufficient to definitively rule out any of them. The significant variation in the $\dot{P}$ across different observations is unlikely to be caused by the intrinsic spin-down of a magnetar. This suggests that either the candidates are due to stochastic processes, or the FRB source resides in a binary system, where the apparent period derivative variations are due to Doppler modulation. Despite the large sample size and the detected short-timescale periodic signals, a definitive period could not be confirmed, indicating that the link between the radiation characteristics of the FRB and spin periodicity may be weak.
  2. 2.
    Long-Timescale Periodicity Analysis: Through epoch folding ($3.9\sigma$) and Lomb-Scargle periodogram ($3.5\sigma$) techniques, we detected a potential $\sim$68-day activity cycle. While the significance is limited by uneven observational coverage, if this periodicity is confirmed, it would place FRB 20240114A among the growing class of periodically active repeaters, possibly indicating binary orbital modulation, precessional effects, or other factors. The lack of robust detection highlights the need for more uniform future observations to constrain long-timescale periodicity in repeating FRBs.
  3. 3.
    Periodicity of burst clusters and QPOs within bursts: We conducted a comprehensive periodicity analysis on the time series of individual bursts and burst clusters to investigate both periodic signals and QPOs in FRBs. For the bursts analysis, our approach, which combines Bayesian inference with EMD, identified 5 significant QPOs ($>3\sigma$), with frequencies in the few hundred Hz range. These QPOs show a strong correlation between central frequency and width (Pearson $R=0.97$), suggesting that the mechanism behind the QPOs may be relatively stable. For the burst clusters analysis, autocorrelation function analysis of burst clusters identified several significant periodic candidates, ranging from a few milliseconds to several tens of milliseconds, with transient coherence over approximately 3–4 cycles. This suggests that oscillation modes are temporarily excited between bursts, providing insight into the magnetospheric plasma dynamics. Collectively, these findings highlight the presence of organized structures both within and between bursts, offering valuable constraints on the physical conditions of the emission region. Based on the relationship between the timescale of pulse microstructure and the rotational period of the source as given in Kramer et al. (2024), we suggest that the possible rotational period of the FRB source is between 2 s and 76 s, although we did not find this period in the data.
FRB 20240114A is currently one of the most active known repeating fast radio bursts, and the results of our periodicity analysis offer important statistical constraints on its physical properties. Through multi-timescale searches for periodic signals and QPOs, we have provided observational clues that may shed light on the source’s rotational characteristics, precession behavior, and dynamic processes occurring both within and between bursts. While these preliminary findings offer a useful starting point, these results clarify the bias of previous studies that detected occasional QPOs believed to be potentially related to the spin of compact objects. As data coverage expands, we hope to refine our understanding of the periodicity and enhance our insights into the connection between fast radio bursts, their rotational properties, and potential precession effects.
acknowledgments
This work made use of the data from FAST FRB Key Science Project. This work is supported by the National Natural Science Foundation of China (NSFC) under grant numbers 12588202, 12303042, 12041303, 12421003, 12233002, 12041306, 12403100, W2442001, 12203045, 12447115, U2031117 and the National SKA Program of China (2020SKA0120100, 2020SKA0120200, 2022SKA0130104). Di Li is a New Cornerstone investigator. Pei Wang acknowledges support from the Youth Innovation Promotion Association CAS (id. 2021055), CAS Youth Interdisciplinary Team. Di Li is supported by the International Partnership Program of Chinese Academy of Sciences, Program No.114A11KYSB20210010, National Key R&D Program of China No. 2023YFE0110500, QN2023061004L. Yi Feng is supported by the Leading Innovation and Entrepreneurship Team of Zhejiang Province of China grant No. 2023R01008, and by Key R&D Program of Zhejiang grant No. 2024SSYS0012. Yongfeng Huang also acknowledges the support from the Xinjiang Tianchi Program. Qin Wu is supported by the China Postdoctoral Science Foundation (CPSF) under grant numbers GZB20240308, 2025T180875. C.W.T is supported by CAS project No. JZHKYPT-2021-06. This research is also supported by the CAS Project for Young Scientists in Basic Research, YSBR-063.
Appendix A Search Method for Periodic Signals and QPOs in bursts
To reliably identify periodic signals and quasi-periodic oscillations (QPOs) in the bursts, we adopt the method proposed by Huppenkothen et al. (2013), which is based on Bayesian inference. The procedure begins by computing the periodogram of the burst time series using the Fast Fourier Transform (FFT). A simpler model, referred to as the null hypothesis model, is then selected, and the periodogram is fit using maximum likelihood estimation (MLE) to obtain the a maximum posteriori (MAP) parameters under this null hypothesis. The analysis is then performed separately for searching periodic signals and QPOs.
For searching periodic signals, the residual at each frequency is computed as:
$R_{j}=\frac{2I_{j}}{S_{j}},$ (A1)
where $I_{j}$ and $S_{j}$ are the values of the periodogram and the null hypothesis model at frequency $f_{j}$, respectively. The maximum residual, $\max(R_{j})$, is identified as a potential periodic signal. The likelihood under the null hypothesis model is calculated as:
$p\left(\bm{I}|\hat{\theta}_{\rm MAP}^{0},H_{0}\right)=\prod\limits_{i=1}^{n}p\left(I_{j}|S_{j}\right),$ (A2)
where $p(I_{j}|S_{j})$ represents the probability density function (PDF) of the periodogram at frequency $f_{j}$ given the model $S_{j}$. The posterior distribution of the null hypothesis model parameters is computed by combining this likelihood with the prior distribution. Markov Chain Monte Carlo (MCMC) sampling is used to obtain a large number of simulated periodograms based on the posterior distribution. The simulated periodograms are then fit with the null hypothesis model, and the maximum residuals for each simulated periodogram are calculated. The $P$-value is determined by comparing the maximum residual observed in the actual burst periodogram with those from the simulated ones. If the $P$-value is below a predefined significance level, the outlier is considered statistically significant and warrants further investigation.
For searching QPOs, the periodogram is fitted with a more complex nested model, referred to as the alternative hypothesis model, to obtain the MAP parameters under this hypothesis. The likelihood ratio is computed by comparing the likelihoods of the periodogram given both the null and alternative hypothesis models:
$T_{\rm LRT}^{\rm obs}=-2\log\frac{p\left(\bm{I}|\hat{\theta}_{\rm MAP}^{0},H_{0}\right)}{p\left(\bm{I}|\hat{\theta}_{\rm MAP}^{1},H_{1}\right)}.$ (A3)
MCMC sampling is also used here to generate simulated periodograms based on the null hypothesis model. The likelihood ratios are calculated for each simulated periodogram, and the $P$-value is determined by comparing the observed likelihood ratio with the sample distribution obtained from the simulations.
The PDF at each frequency in the periodogram is crucial for constructing the likelihood. Many researchers assume that the signal’s power spectral model is multiplied by white noise (Timmer & König, 1995; Vaughan, 2010; Huppenkothen et al., 2013; Li et al., 2022; Xiao et al., 2022), which conflicts with the use of multiple additive models to fit the periodogram. To resolve this issue, we propose a new periodogram PDF based on a time-domain additive model, which more accurately reflects the superposition of multiple components. This approach aligns with Guidorzi (2011), who reached similar conclusions. We first review the periodogram PDF under the frequency-domain multiplication model assumption and then present the PDF under the time-domain additive model assumption.
In the frequency-domain multiplication model, consider a Gaussian white noise $w(t)$ with N time samples, mean $\mu$, and standard deviation $\sigma$. As is well known, its power spectrum at $f>0$ follows ${\rm PDS}(f)=\sigma^{2}$. The periodogram is used to estimate the power spectrum, which can be expressed as
${\rm Per_{W}}(f)=|W(f)|^{2}=\left|\frac{1}{N}\sum_{t=0}^{N-1}w(t)e^{-j2\pi ft}\right|^{2},$ (A4)
where $W(f)$ is the Fourier transform of $w(t)$. It is known that ${\rm Per_{w}}(f)=\frac{1}{2}\sigma^{2}\chi_{2}^{2}$ (Timmer & König, 1995), where $\chi_{2}^{2}$ is a random variable following a chi-squared distribution with 2 degrees of freedom. Therefore, the PDF for the power value ${\rm Per_{w}}(f)=y\ (y>0)$ is
${\rm PDF_{W}}(y;f)=\frac{1}{\sigma^{2}}e^{-\frac{y}{\sigma^{2}}},\quad y>0.$ (A5)
This expression is also equivalent to an exponential distribution with parameter $\lambda=\frac{1}{\sigma^{2}}$. It is considered that for linear processes, the total power spectrum is obtained by multiplying the power spectrum ${\rm PDS_{s}}(f)$ representing the process with the white noise power spectrum ${\rm PDS_{w}}(f)$ (Timmer & König, 1995), i.e., the total power spectrum is ${\rm PDS_{M}}(f)={\rm PDS_{s}}(f)\times{\rm PDS_{w}}(f)$. Therefore, for $f>0$, the probability density of the total periodogram ${\rm Per}(f)=y\ (y>0)$ is
${\rm PDF_{M}}(y;f)=\frac{1}{{\rm PDS}(f)}e^{-\frac{y}{{\rm PDS}(f)}},\quad y>0.$ (A6)
This is consistent with the methods used by Vaughan (2010), Li et al. (2022) and Xiao et al. (2022).
In the time-domain addition model, consider a signal $x(t)$ with N time samples, which consists of a deterministic pulse signal $s(t)$ and a Gaussian white noise signal $w(t)$ with mean $\mu$ and standard deviation $\sigma$, i.e.:
$x(t)=s(t)+w(t).$ (A7)
The Fourie transform of the signal $x(t)$ is
$X(f)=S(f)+W(f),$ (A8)
where $S(f)$ is the Fourie transform of the signal $s(t)$, and the value is deterministic. $W(f)$ is the Fourie transform of the noise $w(t)$, and the value is a complex Gaussian random variable (Timmer & König, 1995), i.e., $W(f)=\mathcal{N}(0,\frac{1}{2}\sigma^{2})+j\mathcal{N}(0,\frac{1}{2}\sigma^{2})$. The total periodogram ${\rm Per_{A}}(f)\equiv|X(f)|^{2}$, and by substituting $X(f)=S(f)+W(f)$, we get
${\rm Per_{A}}(f)=\left|S(f)+W(f)\right|^{2}=\left[{\rm Re}(S(f))+{\rm Re}(W(f))\right]^{2}+\left[{\rm Imag}(S(f))+{\rm Imag}(W(f))\right]^{2}.$ (A9)
It is clear that ${\rm Re}(S(f))+{\rm Re}(W(f))\sim\mathcal{N}({\rm Re}(S(f)),\frac{1}{2}\sigma^{2})$, and ${\rm Imag}(S(f))+{\rm Imag}(W(f))\sim\mathcal{N}({\rm Imag}(S(f)),\frac{1}{2}\sigma^{2})$. Therefore, the power value ${\rm Per_{A}}(f)=\frac{\sigma^{2}}{2}\chi_{2}^{2}\left(\frac{2|S(f)|^{2}}{\sigma^{2}}\right)$, where $\chi_{2}^{2}\left(\frac{2|S(f)|^{2}}{\sigma^{2}}\right)$ is a non-central chi-squared random variable with 2 degrees of freedom and the non-centrality parameter $\lambda(f)=\frac{2|S(f)|^{2}}{\sigma^{2}}$. The PDF for the power value ${\rm Per}(f)=y\ (y>0)$ is
${\rm PDF_{A}}(y;f)=\frac{1}{\sigma^{2}}e^{-\frac{y+|S(f)|^{2}}{\sigma^{2}}}I_{0}\left(\sqrt{\frac{4y|S(f)|^{2}}{\sigma^{4}}}\right),\quad y>0,$ (A10)
where $I_{0}(z)=\sum_{n=0}^{\infty}\frac{1}{(n!)^{2}}\left(\frac{z}{2}\right)^{2n}$ is the modified Bessel function of the first kind, zeroth order. Note that when $|S(f)|^{2}=0$ in equation (A10), the equation degenerates into the result of pure white noise, as shown in equation (A5).
Our search strategy follows that of Huppenkothen et al. (2013) with the following key adjustments: the null hypothesis model consists of a Gaussian model and a constant function model. The Gaussian model fits the low-frequency component of the periodogram, while the constant function model fits the baseline of the periodogram. We found that the Gaussian model provides the best fit for the low-frequency component, and additional Gaussian models are added to fit narrower features in the periodogram. This results in a nested model that serves as the alternative hypothesis. When constructing the likelihood and simulating the periodogram, we use the time-domain additive model PDF (equation A10) rather than the frequency-domain multiplication model PDF (equation A6). After the initial candidate is identified, the periodogram is reviewed manually to determine if additional Gaussian models are needed. If the periodogram cannot be fit with at most five Gaussian functions plus one constant function (with a reduced $\chi^{2}>1.5$), the candidate is discarded.
While the $P$-values reported in this work already account for multiple testing across frequencies in each individual periodogram, we do not apply additional corrections for multiple bursts since we focus on assessing significance within each burst individually. This conservative approach avoids over-correction, especially since previously published bursts are not fully included in our analysis.
Appendix B QPO fitting parameters for individual bursts
Table 2: QPO fitting parameters for individual bursts. The table lists the model components (Model), their associated amplitudes (Height), central frequencies (CF), full width at half maximum (FWHM), quality factors (Q), and constant (C) for the QPO in each individual burst. The MJD in each row indicates the start time of the time series for that particular burst.
No. MJD DM Model Height CF FWHM Q C
From 60355 $\rm pc/cm^{3}$ Hz Hz ×$\times 10^{-3}$
1 20.182604131 528.5 G0 6.337 $\pm$ 0.046 4.45 $\pm$ 0.59 85.02 $\pm$ 0.63 0.0523 $\pm$ 0.0069 -
G1 0.1865 $\pm$ 0.0037 266.85 $\pm$ 0.68 97.0 $\pm$ 1.3 2.750 $\pm$ 0.037 -
G2 0.0053 $\pm$ 0.0011 502.1 $\pm$ 6.2 70 $\pm$ 14 7.1 $\pm$ 1.4 -
Const - - - - $3.0\pm 4.6$
2 20.200749305 527.8 G0 160.4 $\pm$ 2.3 0.8 $\pm$ 1.4 115.0 $\pm$ 1.4 0.007 $\pm$ 0.012 -
G1 14.22 $\pm$ 0.18 277.90 $\pm$ 0.61 124.7 $\pm$ 1.3 2.228 $\pm$ 0.024 -
G2 0.592 $\pm$ 0.029 501.8 $\pm$ 8.3 212 $\pm$ 12 2.36 $\pm$ 0.14 -
Const - - - - $2.7\pm 2.9$
3 21.143877879 528.7 G0 36.3 $\pm$ 0.4 5.8 $\pm$ 1.1 118.5 $\pm$ 1.1 0.0487 $\pm$ 0.0094 -
G1 0.946 $\pm$ 0.021 327.94 $\pm$ 0.93 117.4 $\pm$ 1.7 2.793 $\pm$ 0.041 -
G2 0.0368 $\pm$ 0.0047 579.8 $\pm$ 5.6 105 $\pm$ 10 5.50 $\pm$ 0.55 -
Const - - - - $2.2\pm 1.2$
4 26.127338551 527.3 G0 9.04 $\pm$ 0.28 0.0 $\pm$ 2.7 253.2 $\pm$ 5.9 $0$$\pm$ 0.011 -
G1 0.212 $\pm$ 0.012 586.6 $\pm$ 4.4 212.0 $\pm$ 8.6 2.77 $\pm$ 0.11 -
Const - - - - $2.2\pm 2.0$
5 114.836935924 528.3 G0 6.02 $\pm$ 0.14 0.0 $\pm$ 2.9 176.0 $\pm$ 2.7 $0$$\pm$ 0.016 -
G1 0.0980 $\pm$ 0.0062 468.1 $\pm$ 3.9 159.8 $\pm$ 7.1 2.93 $\pm$ 0.13 -
Const - - - - $2.3\pm 1.4$
Appendix C Quasi-Periodic Signals in Burst Time Series
Figure 9: Detections of QPOs in the burst time series of FRB 20240114A. Each candidate is displayed in a panel of four subplots, with its identifier labeled in the top-left subplot. For each candidate, the left column shows the original burst time series, the EMD result (black curve), and the fitted model from equation (8) (pink curve), while the right column displays the corresponding periodogram, where dashed lines indicate the best-fit components and the central frequencies of QPOs.
Figure 10: Candidate periodic signals identified from the burst time series using the ACF method, each showing periodic indications with some significance, although not exceeding $3\sigma$. Each candidate consists of four subplots: the top-left subplot shows the burst time series, the bottom-left subplot shows the corresponding dynamic spectrum, the top-right subplot shows the ACF of the burst time series, and the bottom-right subplot shows the simulated statistic used to calculate the probability of generating this ACF pattern. The red line in the burst time series represents the periodic intervals determined by the ACF analysis, while the red line in the simulated statistic corresponds to the observed value.